OpenStudy (waheguru):

how to isolate h 2pir^2+2pirh=v???????

jimthompson5910 (jim_thompson5910):

\[\Large 2\pi r^2 + 2\pi r h = v\] \[\Large 2\pi r h = v-2\pi r^2\] \[\Large h = \frac{v-2\pi r^2}{2\pi r}\]